OpenStudy (anonymous):

Solve the inequality. 3/10 greater than or equal to k - 3/5

OpenStudy (anonymous):

\[\frac{ 3 }{ 10 }\ge k-\frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

yeah thats it

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

now isolate (solve) for k by adding \[\frac{ 3 }{ 5 }\] to each side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you get \[\frac{ 3 }{ 10 }+ \frac{ 3 }{ 5 }\ge k\]

OpenStudy (anonymous):

now add the fractions... what to we get?

OpenStudy (anonymous):

3/10

OpenStudy (anonymous):

you don't know how to add fractions?

OpenStudy (anonymous):

find the common denominator between 5 and 10, which is 10... you must change then denominators ( bottom number of the fraction) to equal eachother in order to add them).

OpenStudy (anonymous):

so let's change 3/5 to 6/10 so we can add it to 3/10

OpenStudy (anonymous):

3/5=6/10

OpenStudy (anonymous):

\[\frac{ 3 }{ 10 }+ \frac{ 6 }{ 10}\ge k\]